# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v3

@@ IV.3.1 Four-particle symmetric cumulants
Next, we compare the results from our simulations with the experimental measurements from RHIC. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [44, 42, 37, 45, 46, 47, 48], while ${\rm AC}_{2,1}(m,n)$ have been computed in some transport models [46, 47]. Fig. [5] shows comparisons between the theory calculations for the symmetric cumulants ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ multiplied by $\langle N_{part}\rangle$ and the same measured by the STAR Collaboration [73]. The factor of $\langle N_{part}\rangle$ is multiplied to scale out the trivial dilution of correlation with the increase in the number of multiplets. We are able to reproduce the qualitative variations with centrality. The magnitudes of $\langle N_{part}\rangle{\rm SC}(2,3)$ and $\langle N_{part}\rangle{\rm SC}(2,4)$ increase from central to mid-peripheral events and then again decreases for very peripheral events. Here, we do not aim to reproduce the experimental data because our aim is to check the sensitivity of these observables to the transport properties of the system created in these collisions. The comparison in Fig.[5] reflects that the correlation in initial spatial eccentricities and the following hydrodynamic evolution can capture the correlation among flow harmonics of different orders. Negative values of ${\rm SC}(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as ${\rm SC}(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. Symmetric cumulants do not have contributions from non-flow effects, where non-flow refers to azimuthal correlations not related to the reaction plane orientation, like those from resonances, jets, quantum statistics, etc. This is verified by computing these observables for the HIJING model, which includes only non-flow physics, for which these are consistently zero [35]. The model captures the variation of ${\rm SC}(m,n)$ with the centrality. In central collisions, the model underestimates ${\rm SC}(2,4)$, possibly due to inadequate fluctuations in the initial conditions or some shortcomings in the transport model since the nonlinear response in $v_{4}$ and $\varepsilon_{4}$ is sensitive to these effects. However, ${\rm SC}(2,3)$ is better aligned with the data because the $v_{2}-v_{3}$ correlations mainly arise from the initial state $\varepsilon_{2}-\varepsilon_{3}$ correlations due to linear response in $v_{2}$ and $v_{3}$.
@@
Figure 5: (Normalized) Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [73] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive.
@@
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq.[5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity (Eq.[11]) evaluated using the Monte Carlo Glauber model, are presented in Fig.[5]. We observe that ${\rm NSC}(2,3)$ shows little sensitivity to shear and bulk viscosities. However, ${\rm NSC}(2,4)$ is more sensitive to both shear and bulk viscosities than $v_{2}\{4\}$. This is due to the fact that $v_{4}$ is the superposition of a nonlinear term that is proportional to $(v_{2})^{2}$ and a linear term $v_{4L}$. This can be expressed as:
$v_{4}=v_{4L}+\chi_{4}(v_{2})^{2},$
where $\chi_{4}$ is the non-linear response coefficient. The term $v_{4L}$ is damped more strongly than the nonlinear term, as noted in Reference [74]. Additionally, $v_{4L}$ is less correlated with $v_{2}$. Consequently, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$ also increases.
@@ IV.3.2 Six-particle asymmetric cumulants
${\rm AC}_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [50]. In Fig.[6], we show the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$ to scale out the trivial dilution of correlation with the increase in the number of multiplets, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compute these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities. As discussed earlier, the linear component in $v_{4}$ is damped more significantly than the nonlinear term and shows a weaker correlation with $v_{2}$. Thus, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$ is also enhanced.
@@
Figure 6: (Normalized) Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive.
@@
In Fig.[6], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq.[8] and the corresponding equation for the initial state eccentricities (Eq.[11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}_{2,1}(m,n)$ in the final state to be equal to the ${\rm NAC}_{2,1}(m,n)$ in the initial state. Fig. [6] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. The correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is weaker than the correlation observed between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ comes not only from the linear response of the system to $\varepsilon_{4}$, but also from a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NAC}(2,4)$ in the final state and the initial state increases. This is likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$.
@@
Currently, STAR has not yet measured asymmetric cumulants. We conducted a comparison between our findings and the Pb+Pb 5.02 TeV results from ALICE (See Appendix Fig.[7] and Fig.[8]). The centrality dependence of the observables we discussed in this paper shows similarities between ALICE and STAR data. The ${\rm SC}(m,n)$ values are about an order of magnitude larger at ALICE than at STAR. Similarly, the ${\rm AC}_{21}(m,n)$ values are 3 to 5 times larger at ALICE compared to STAR. However, the similarities in the magnitude and centrality dependence of the normalized cumulants are noteworthy. This comparison indicates minimal energy dependence for the normalized cumulants.
IV.4 Conclusions
@@
We have presented results for multi-particle correlation functions in heavy-ion collisions at top RHIC energy using a hybrid framework based on the Monte-Carlo Glauber model, MUSIC viscous hydrodynamics simulations, and the UrQMD hadronic cascade.
@@
First, we adjusted the free parameters, such as shear and bulk viscosities, to describe particle multiplicities, mean transverse momentum, and anisotropic flow. After that, we discussed the impact of shear and bulk viscosities, resonance decay, and hadronic interactions on traditional and novel observables.
@@
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are very sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. However, this magnified sensitivity can be attributed to the properties of anisotropic flow, particularly as the cumulants incorporate higher powers of flow. We then checked the sensitivities of the normalized cumulants. We found that ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$ are insensitive to both the hydro model parameters (shear and bulk viscosity) and late-stage hadronic interactions. This strongly supports their effectiveness as reliable tools for constraining the initial state of the system’s evolution. Conversely, ${\rm NSC}(2,4)$ and ${\rm NAC}_{2,1}(2,4)$ show considerable sensitivity to these stages, allowing them to effectively constrain hydrodynamic and transport models.
A single parameter set of transport coefficients is unable to explain $SC(2,3)$ and $SC(2,4)$ simultaneously.
@@
We observed anti-correlation between event-by-event fluctuations of $v_{2}$ and $v_{3}$, while the event-by-event fluctuations of $v_{2}$ and $v_{4}$ are found to be positively correlated. The observed anti-correlation between $v_{2}$ and $v_{3}$ appears to be described by the initial-stage anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which supports the idea of linearity between $\varepsilon_{n}$ and $v_{n}$ [23]. The hydrodynamic response suggests that the final state fluctuation originates from the initial state. However, including the nonlinear hydrodynamic response of the medium is necessary to reproduce the measured correlation between $v_{2}$ and $v_{4}$, as the initial-stage linear correlation alone is insufficient.
@@
The values of ${\rm SC}(m,n)$ and ${\rm AC}_{21}(m,n)$ at ALICE are larger—by factors ranging from 3 to 8—compared to those at STAR. However, it is important to note the similarities in the magnitude and centrality dependence of the normalized cumulants. This comparison indicates that there is minimal energy dependence for the normalized cumulants.
IV.5 Acknowledgements
We would like to acknowledge Tribhuban Parida for fruitful discussions and providing us the computational setup.

### v4

@@ IV.3.1 Four-particle symmetric cumulants
Next, we compare the results from our simulations with the experimental measurements from RHIC. Fig. [5] shows comparisons between the theory calculations for the symmetric cumulants ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ multiplied by $\langle N_{part}\rangle$ and the same measured by the STAR Collaboration [37]. The factor of $\langle N_{part}\rangle$ is multiplied to scale out the trivial dilution of correlation with the increase in the number of multiplets. The magnitudes of $\langle N_{part}\rangle{\rm SC}(2,3)$ and $\langle N_{part}\rangle{\rm SC}(2,4)$ increase from central to mid-peripheral events and then again decreases for very peripheral events. Here, we do not aim to reproduce the experimental data because our aim is to check the sensitivity of these observables to the transport properties of the system created in these collisions. The comparison in Fig. [5] reflects that the correlation in initial spatial eccentricities and the following hydrodynamic evolution can capture the correlation among flow harmonics of different orders. Negative values of ${\rm SC}(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as ${\rm SC}(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. Symmetric cumulants do not have contributions from non-flow effects, where non-flow refers to azimuthal correlations not related to the reaction plane orientation, like those from resonances, jets, quantum statistics, etc. This is verified by computing these observables for the HIJING model, which includes only non-flow physics, for which these are consistently zero [35]. The model reproduces the qualitative variation of ${\rm SC}(m,n)$ with the centrality. However, a single set of transport coefficients cannot simultaneously explain both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$.
@@
Figure 5: (Normalized) Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [37] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive. Note that we have shown the Glauber results for ${\rm NSC}(2,4)$ to highlight that in non-central collisions $v_{4}$ is mainly driven by $\varepsilon_{2}^{2}$ and not $\varepsilon_{4}$.
@@
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq. [5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity, $\varepsilon_{n}$ (Eq. [11]), evaluated using the Monte Carlo Glauber model, are presented in Fig. [5].
@@
We observe that ${\rm NSC}(2,3)$ shows little sensitivity to shear and bulk viscosities. However, ${\rm NSC}(2,4)$ is more sensitive to both shear and bulk viscosities than $v_{2}\{4\}$. This is due to the fact that $v_{4}$ is the superposition of a linear term and a nonlinear term. This can be expressed as [72]:
$v_{4}e^{-i4\psi_{4}}=w_{4}e^{-i4\phi_{4}}+w_{4(22)}e^{-i4\phi_{2}},$ (12)
where $w_{4}$ represents the linear response while $w_{4(22)}$ describes the nonlinear response. Notably, the linear term, $w_{4}$, experiences greater damping compared to the nonlinear term, $w_{4(22)}$, as noted in Reference [72]. Furthermore, $w_{4}$ demonstrates a weaker correlation with $v_{2}$. Consequently, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$ also increases  [^1] We thank the anonymous referee for their insightful comments and suggestions, particularly the detailed explanation that helped us expand the context of this result..
@@ IV.3.2 Six-particle asymmetric cumulants
In Fig. [6], we show the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. As for symmetric cumulants, negative values of ${\rm AC}_{2,1}(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as ${\rm AC}_{2,1}(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compare these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities.
@@
Figure 6: (Normalized) Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive. Note that we have shown the Glauber results for ${\rm NAC}_{2,1}(2,4)$ to highlight that in non-central collisions $v_{4}$ is mainly driven by $\varepsilon_{2}^{2}$ and not $\varepsilon_{4}$.
@@
In Fig. [6], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq. [8] and the corresponding equation for the initial state eccentricities (Eq. [11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}_{2,1}(m,n)$ in the final state to be equal to the ${\rm NAC}_{2,1}(m,n)$ in the initial state. Fig. [6] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. Similar to the reasons mentioned for ${\rm NSC}(2,4)$, as collisions become more peripheral, the difference in ${\rm NAC}_{2,1}(2,4)$ between the initial and the final states increases. We compare these observables for the three sets of ensembles mentioned in Table  2 and observe that, while ${\rm NAC}_{2,1}(2,3)$ remains unaffected, ${\rm NAC}_{2,1}(2,4)$ is notably enhanced by the influence of shear and bulk viscosities. As discussed earlier, the linear component in $v_{4}$ is damped more significantly than the nonlinear term and shows a weaker correlation with $v_{2}$. Consequently, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$, and thus ${\rm NAC}_{2,1}(2,4)$, is also enhanced.
@@
Currently, STAR has not yet measured asymmetric cumulants. We conducted a comparison between our findings and the Pb+Pb 5.02 TeV results from ALICE (See Fig. 7 in Appendix VII). The centrality dependence of the observables we discussed in this paper shows similarities between ALICE and STAR data. The ${\rm SC}(m,n)$ values are about an order of magnitude larger at ALICE than at STAR. Similarly, the ${\rm AC}_{2,1}(m,n)$ values are 3 to 5 times larger at ALICE compared to STAR. However, the similarities in the magnitude and centrality dependence of the normalized cumulants are noteworthy. This comparison indicates minimal energy dependence for the normalized cumulants.
V Conclusions
@@
We have presented results for multi-particle correlation functions in heavy-ion collisions at top RHIC energy using a hybrid framework based on the Monte Carlo Glauber model, MUSIC viscous hydrodynamics simulations, and the UrQMD hadronic cascade.
@@
First, we tuned the free parameters, such as shear and bulk viscosities, to describe particle multiplicities, mean transverse momentum, and anisotropic flow. Subsequently, we discussed the impact of shear and bulk viscosities, resonance decay, and hadronic interactions on traditional and novel observables.
@@
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are highly sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. However, this magnified sensitivity can be attributed to the properties of anisotropic flow, particularly as the cumulants incorporate higher powers of flow. We then examined the sensitivities of the normalized cumulants. We found that ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$ are insensitive to both the hydro model parameters (shear and bulk viscosity) and late-stage hadronic interactions. This strongly supports their effectiveness as reliable tools for constraining the initial state of the system’s evolution. Conversely, ${\rm NSC}(2,4)$ and ${\rm NAC}_{2,1}(2,4)$ show considerable sensitivity to these stages, allowing them to effectively constrain hydrodynamic and transport models.
@@
We observed anti-correlation between event-by-event fluctuations of $v_{2}$ and $v_{3}$, while the event-by-event fluctuations of $v_{2}$ and $v_{4}$ are found to be positively correlated. The observed anti-correlation between $v_{2}$ and $v_{3}$ appears to be described by the initial-stage anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which supports the idea of linearity between $\varepsilon_{n}$ and $v_{n}$ [23]. The hydrodynamic response suggests that the final state fluctuation originates from the initial state. However, it is important to consider the nonlinear hydrodynamic response of the medium to explain the measured correlation between $v_{2}$ and $v_{4}$, as the initial-stage linear correlation alone is insufficient.
@@
The values of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ at ALICE are larger—by factors ranging from 3 to 8—compared to those at STAR. However, it is important to note the similarities in the magnitude and centrality dependence of the normalized cumulants. This comparison indicates that there is minimal energy dependence for the normalized cumulants.
VI Acknowledgements
We would like to acknowledge Tribhuban Parida for fruitful discussions and providing us the computational setup. We would like to express our sincere gratitude to the anonymous referee for their invaluable insights and constructive feedback, especially for the thorough and thoughtful explanation they provided regarding one of our key results.